# ECORE: Energy-Conscious Optimized Routing for Deep Learning Models at the Edge

### v2

@@ Abstract.
Edge computing enables data processing closer to the source, significantly reducing latency—an essential requirement for real-time vision-based analytics such as object detection in surveillance and smart city environments. However, these tasks place substantial demands on resource-constrained edge devices, making the joint optimization of energy consumption and detection accuracy critical. To address this challenge, we propose ECORE , a framework that integrates multiple dynamic routing strategies—including estimation-based techniques and a greedy selection algorithm—to direct image processing requests to the most suitable edge device–model pair. ECORE dynamically balances energy efficiency and detection performance based on object characteristics. We evaluate our approach through extensive experiments on real-world datasets, comparing the proposed routers against widely used baseline techniques. The evaluation leverages established object detection models (YOLO, SSD, EfficientDet) and diverse edge platforms, including Jetson Orin Nano, Raspberry Pi 4 and 5, and TPU accelerators. Results demonstrate that our proposed context-aware routing strategies can reduce energy consumption and latency by 45% and 49%, respectively, while incurring only a 2% loss in detection accuracy compared to accuracy-centric methods.
Edge computing, Dynamic Routing, Object Detection Models, Energy Consumption, Accuracy.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; May 11-14, 2026; Saint-Malo, Franceisbn: XXX-X-XXXX-XXXX-X/2025/06
1. Introduction
The rapid growth of the Internet of Things (IoT) and smart applications in recent years has led to a massive increase in data generated by distributed sensing devices. According to a forecast by the International Data Corporation (IDC) (Reinsel, 2019), there will be 55.9 billion IoT devices generating approximately 79.4 zettabytes (ZB) of data by 2025. Continuously transmitting this volume of data to centralized data centers or cloud platforms for processing is often infeasible due to several limitations, including network bandwidth constraints, increased latency, and potential privacy and security concerns. These challenges are particularly significant in applications that require real-time decision-making or operate in bandwidth-limited environments. As a response, edge computing has emerged as a viable solution, enabling data processing to occur closer to the source–on or near the IoT devices using resource constrained single board computers such as Jetson, Raspberry Pi, and similar edge platforms. By offloading computation to the edge, this paradigm reduces communication overhead, lowers response times, and enhances the scalability and autonomy of intelligent systems.
@@ 1. Introduction
Object detection, a key task in computer vision-based analytics, has seen widespread adoption across various domains including agriculture, sports analytics, transportation, healthcare, and smart cities. Recent advancements in machine learning and artificial intelligence (AI) have significantly improved the performance of object detection models. In particular, deep learning-based models such as You Only Look Once (YOLO) (Redmon et al., 2016), Single shot multibox detector (SSD) (Liu et al., 2016) and EfficientDet (Tan et al., 2020) have become increasingly prominent due to their high detection accuracy and real time capabilities. These models are often implemented using convolutional neural network (CNN) architectures and come in multiple variants to accommodate different computational budgets. For instance, the YOLOv8 (Jocher et al., 2023) family includes nano, small, medium, and large versions each reflecting trade-offs between model size, speed, and accuracy.
@@
The dynamic nature of object detection in real-world environments presents both a challenge and an opportunity. Visual scenes vary widely in complexity—from sparsely populated streets at night to dense daytime urban intersections—leading to fluctuating computational demands for object detection. In such contexts, the heterogeneity of edge resources, including both hardware devices and detection models, offers an opportunity for intelligent allocation of requests to device–model pairs.
Our recent findings show that detection models exhibit notably different energy, latency, and accuracy profiles when deployed across various edge devices (Alqahtani et al., 2024). By leveraging this hardware–model heterogeneity, systems can achieve substantial energy savings without compromising detection accuracy or latency.
@@
Energy efficiency is particularly critical on edge devices operating under strict power constraints, such as those relying on batteries or local energy harvesting methods like solar panels. Yet in safety-critical applications like surveillance and traffic monitoring, accuracy remains non-negotiable. This creates the central challenge addressed in this study: “how can we minimize energy usage without degrading detection quality?
@@
The solution lies in adaptive, context-aware routing strategies that dynamically select the optimal device–model pair based on scene complexity. Simpler scenes allow for lightweight processing, while complex environments require more accurate–and consequently more energy-intensive–models. To address this, we propose the ECORE framework, which introduces a set of dynamic, novel, intelligent routing algorithms to redirect an image processing task to the most appropriate device-model pair. The goal is to reduce energy consumption while maintaining high detection accuracy.
@@
ECORE’s routers operate by applying lightweight complexity estimation techniques during the pre-processing stage to infer the number of objects in each incoming image. Based on this estimation, the router selects and forwards the request to an optimal model–device pair.
The main contributions of this work are as follows:
  1. (1)
@@
    We propose a complete system architecture for a framework called ECORE , optimized routing of image processing tasks to a group of heterogeneous edge devices running diverse object detection models, coordinated by a central gateway responsible for intelligent routing decisions.
  2. (2)
@@
    We develop three object count estimation approaches to support intelligent routing decisions in ECORE . The first is an edge detection–based approach, which utilizes the Canny technique to estimate the number of objects in an image. The second is an SSD-based front-end approach, which deploys a lightweight object detection model at the gateway to make decisions. The third is an outputs-based approach, which leverages the object detection results returned from the backbend in the previous processing round and uses this information to inform routing decisions for the subsequent request.
  3. (3)
@@
    We conduct a comprehensive experimental evaluation of our proposed routing approaches by comparing them against both an ideal benchmark–an Oracle Router that knows the exact number of objects in the image–and several baseline techniques, including Random, Round Robin, Lowest Energy Consumption, Lowest Inference Time, and Highest mean Average Precision (mAP). The evaluation is conducted on a real-world testbed using multiple datasets, including the Common Objects in Context (COCO) validation set and a short video, under diverse input scenarios.
@@
The remainder of the paper is organized as follows: Section 2 discusses the motivation behind this work. Section 3 presents the system design, problem formulation, and greedy routing approach. Section 4 describes the datasets, experimental setup, and results. Section 5 reviews related literature. Finally, Section 6 concludes the paper and outlines directions for future work.
2. Motivation
@@
This section highlights the motivation for developing adaptive, context-aware routing strategies in edge-based object detection systems. Figure LABEL:fig:scene_examples illustrates two contrasting scenes captured by an urban surveillance camera at an intersection. The first image depicts numerous pedestrians crossing the intersection, while the second image shows the same location when it is sparsely populated. In the former high density situation, accurate object detection becomes more critical and potentially more computationally demanding, especially in safety sensitive applications such as surveillance, traffic monitoring, and crowd management. Failure to detect or correctly classify objects in this context can compromise system reliability and lead to safety risks. In contrast, the latter image shows a sparsely populated street with minimal environmental activity and few objects of interests. In such low activity scenarios, the computational complexity of achieving reasonable accuracy is reduced, and the risk of critical misdetections is substantially lower. Therefore, energy savings can be achieved by using lighter detection models, particularly when edge devices operate under battery or solar power constraints.
These scenarios underscore the importance of a dynamic routing approach that forwards requests to the optimal edge node based on scene complexity. By estimating the complexity of a scene prior to routing, the system can intelligently choose between high accuracy models for crowded scenes and energy efficient models for sparse scenes. This motivation drives the design of our proposed routing algorithm, which aims to optimize energy consumption without compromising accuracy where it matters most.
@@
To verify our hypothesis, we conducted a preliminary experiment using the COCO validation dataset, which contains 5,000 images. We categorized the images into two groups based on the number of objects, as determined by the ground truth: one group with a single object per image and another with four or more. We then calculated the mAP and energy consumption per image during inference. As shown in Figure LABEL:fig:energy_and_mAP(b), SSD Lite and YOLOv8n exhibit similar accuracy for single-object images, while YOLOv8n significantly outperforms SSD Lite on images with four or more objects—achieving nearly twice the mAP. In terms of energy usage, SSD Lite consumes a consistent amount across both groups, which is approximately 50% lower than YOLOv8n. These results suggest that in less complex scenarios, two different object detection models can achieve similar accuracy, while significant energy savings is possible by using more efficient models without compromising accuracy.
3. System Design and Implementation
@@
This section describes the system architecture of ECORE , designed for energy-efficient and performance-aware routing of object detection tasks in an edge–IoT environment. Figure 2 illustrates the overall design. We assume a scenario where a pool of backend edge nodes hosts a variety of model–device pairs. Each backend device runs a dedicated object detection model and exposes an inference API to process images or video frames captured via distributed cameras or embedded sensors. This modular and distributed architecture ensures scalability and responsiveness, making it suitable for real-time applications in energy-constrained and performance-critical edge computing scenarios.
@@
Cameras or embedded sensors transmit image data or video frames to a central gateway, which performs lightweight preprocessing and analysis. As shown in Figure 2, the gateway first applies one of the proposed object estimation techniques to infer the number of objects in each incoming image. This count is passed to the routing algorithm (Algorithm 1), which uses it to determine the most appropriate model–device pair.
@@
The routing algorithm begins by identifying the corresponding object count group based on the estimated number of objects, using a predefined set of group rules defined by numeric ranges and labels. It then filters the profiling data to retain only the entries that match the identified group. From this subset, it calculates the maximum achievable mAP and derives a minimum acceptable mAP by subtracting a predefined margin. The data is then filtered again to retain only model–device pairs whose mAP falls within this target range. Among these, the pair with the lowest energy consumption is selected. The gateway forwards the request to the selected endpoint, which processes the image and returns the detected objects and their bounding boxes.
Figure 2. System architecture presentation.
3.1. Problem Formulation and Greedy Approach
@@
When running an object detection application in an edge environment, a key challenge lies in routing image requests to the most suitable model-device pair, taking into account trade-offs among accuracy, latency, and energy consumption. Let
=$\mathcal{M}=\{(m_{1},d_{1}),(m_{2},d_{2}),\dots,(m_{n},d_{n})\}$
denote the set of available object detection model-device pairs, where $m_{i}$ represents a specific object detection model and $d_{i}$ denotes the device on which it can be executed. Each pair $(m_{i},d_{i})\in\mathcal{M}$ is profiled in advance, yielding key performance indicators:
  • $\text{mAP}_{i}$: mean Average Precision
  • $t_{i}$: inference time (in milliseconds)
  • $e_{i}$: energy consumption (in milliwatt-hour (mWh))
  • $g_{i}$: object count group (e.g. 0, 1, 2, 3, etc)
@@
=$\mathcal{F}=\{i\in\mathcal{M}_{G}\mid\text{mAP}_{i}\geq\text{mAP}_{\text{max}}-\delta_{\text{mAP}}\}$
@@
[…]
@@
[…]
@@
[…]
@@
This greedy algorithm, as detailed in Algorithm 1, provides a principled balance between performance and efficiency, enabling the system to route requests to energy-efficient configurations with only marginal degradation in accuracy. It is particularly suitable for real-time applications deployed in resource constrained environments, where the router must make fast, intelligent decisions under uncertain and dynamic workloads.
3.2. Optimality of the Routing Algorithm
@@
Theorem 3.1 (Optimality of the Routing Algorithm).
@@
=$\mathcal{F}=\left\{i\in\mathcal{M}_{G}\mid\text{mAP}_{i}\geq\text{mAP}_{\max}-\delta_{\text{mAP}}\right\}$
@@
The routing objective is to minimize energy consumption:
@@
where the feasible set $\mathcal{F}$ includes only those model-device pairs in $\mathcal{M}_{G}$ whose accuracy exceeds the threshold $\text{mAP}_{\max}-\delta_{\text{mAP}}$.
@@
As the energy consumption values $e_{i}$ are known from prior profiling and independent of each other, the greedy algorithm that selects
@@
3.3. Proposed object count estimation approaches
It is crucial to keep the core of the routing algorithm proposed in the previous subsection—particularly the component responsible for estimating the number of objects in an image—lightweight. Otherwise, the overhead introduced by the algorithm could offset the energy savings achieved by the system. Therefore, to estimate the number of objects in incoming image requests and redirect them to the most suitable edge node, we develop and evaluate three estimation approaches:
  1. (1)
    Edge Detection (ED): This approach leverages edge detection which is a fundamental technique in image processing and computer vision used to identify areas of significant intensity or color change, which often correspond to object boundaries. The primary goal is to capture essential structural information. We employ the widely used Canny edge detection algorithm (Canny, 1986), to estimate the number of objects in each image. The resulting object count is then fed into the proposed routing algorithm (Algorithm 1), which selects the optimal model-device pair based on the expected energy–accuracy trade-off governed by $\delta_{\text{mAP}}$. This method is considered lightweight and computationally efficient, though it may yield coarse or approximate estimations.
  2. (2)
@@ 3.3. Proposed object count estimation approaches
  3. (3)
    Output-Based (OB): This approach leverages temporal continuity in image streams—particularly video—by assuming that consecutive frames contain a similar number of objects. Instead of estimating the object count for every image (frame), the system reuses the count obtained from the output of the previous image (frame), as detected by the last selected model–device pair. The same configuration is retained until a change in object count is observed. The routing algorithm begins with a default estimate (e.g., zero) and forwards the first image to an edge node. Once the detection results are returned, the router extracts the actual number of objects and uses it for subsequent routing decisions. If the object count remains stable, the system continues using the same model–device pair; if a change is detected, the routing decision is updated accordingly. This adaptive strategy avoids per-image (frame) estimation, reducing computational overhead and improving energy efficiency. OB is particularly effective for video streams or sequences of images that exhibit consistent object counts or computational complexity, where the number of objects tends to vary gradually over time.
4. Performance Evaluation
4.1. Experimental Setup
4.1.1. Dataset
In this study, we utilize three distinct datasets, two of which are derived from the COCO validation set (Lin et al., 2014), to evaluate the routing algorithm. The first is the Original COCO Validation Dataset, consisting of 5,000 images annotated with ground truth labels, including object classes and bounding boxes spanning 80 object categories. As this dataset reflects natural scene variability, the number of objects per image varies significantly. Figure 3 illustrates the distribution of object counts, highlighting this diversity.

### v3

@@ Abstract
Edge computing enables data processing closer to the source, significantly reducing latency, an essential requirement for real-time vision-based analytics such as object detection in surveillance and smart city environments. However, these tasks place substantial demands on resource-constrained edge devices, making the joint optimization of energy consumption and detection accuracy critical. To address this challenge, we propose ECORE , a framework that integrates multiple dynamic routing strategies, including a novel estimation-based techniques and an innovative greedy selection algorithm, to direct image processing requests to the most suitable edge device–model pair. ECORE dynamically balances energy efficiency and detection performance based on object characteristics. We evaluate our framework through extensive experiments on real-world datasets, comparing against widely used baseline techniques. The evaluation leverages established object detection models (YOLO, SSD, EfficientDet) and diverse edge platforms, including Jetson Orin Nano, Raspberry Pi 4 and 5, and TPU accelerators. Results demonstrate that our proposed context-aware routing strategies can reduce energy consumption and latency by 35% and 49%, respectively, while incurring only a 2% loss in detection accuracy compared to accuracy-centric methods.
I Introduction
@@
Object detection, a core task in computer vision analytics, has been widely adopted across domains such as agriculture, sports analytics, transportation, healthcare, and smart cities. Recent advancements in machine learning and artificial intelligence (AI) have significantly enhanced the performance of object detection models. As a result, deploying these models on edge platforms such as Raspberry Pi and NVIDIA Jetson has become an effective strategy for edge computing, thereby reducing latency, preserving privacy, and enabling scalable, on-device intelligence. In particular, deep learning-based models such as You Only Look Once (YOLO) [1], Single shot multibox detector (SSD) [2] and EfficientDet [3] come in multiple variants to accommodate different computational budgets. For instance, the YOLOv8 [4] family includes nano, small, medium, and large versions each reflecting trade-offs between model size, speed, and accuracy.
@@
The dynamic nature of object detection in real-world environments presents both a challenge and an opportunity. Visual scenes vary widely in complexity, from sparsely populated streets at night to dense daytime urban intersections, leading to fluctuating computational demands for object detection. In such contexts, the heterogeneity of edge resources, including both hardware and detection models, offers an opportunity for intelligent allocation of requests to device–model pairs.
@@
Our recent findings show that detection models exhibit markedly different energy, latency, and accuracy profiles when deployed diverse edge devices [5]. By leveraging this hardware–model diversity, systems can achieve substantial energy savings without compromising detection accuracy or latency. Energy efficiency is particularly critical on edge devices operating under strict power constraints, such as those relying on batteries or local energy harvesting methods like solar panels. Yet in safety-critical applications like surveillance and traffic monitoring, accuracy remains non-negotiable. This creates the central challenge addressed in this study: “how can we minimize energy usage without degrading detection quality?
@@
Our proposed solution lies in adaptive, context-aware routing strategies that dynamically select the optimal device–model pair based on scene complexity. Simpler scenes allow for lightweight processing, while complex environments require more accurate–and consequently more energy-intensive–models. To address this, we propose the ECORE framework, which introduces a set of dynamic and intelligent routing algorithms to redirect an image processing task to the most appropriate device-model pair. The goal is to reduce energy consumption while maintaining high detection accuracy. Note that this routing mechanism should be lightweight, so it does not offset the energy savings, since any added overhead increases the overall image processing pipeline cost.
@@
ECORE’s routers use lightweight estimation techniques during the pre-processing stage to infer the number of objects in each incoming image as an indicator of scene complexity. To the best of our knowledge, this is the first work to address the routing problem through the lens of scene complexity. Based on this estimate, the router selects and forwards the request to an optimal model–device pair. The main contributions of this work are as follows:
  1. 1.
@@
    A first-of-its-kind system architecture for a framework called ECORE , providing optimized routing of image processing tasks across heterogeneous edge devices running diverse object detection models, coordinated by a central gateway responsible for routing decisions.
  2. 2.
@@
    Three novel object count estimation approaches to support intelligent routing decisions in ECORE . The first is a light-weight edge detection–based approach, which utilizes the Canny technique to estimate the number of objects in an image. The second is an SSD-based front-end approach, which deploys a minimal object detection model at the gateway to make decisions. The third is low-overhead approach approach, which leverages the object detection outputs returned from the backbend in the previous processing round and uses this information to inform routing decisions for the subsequent request.
  3. 3.
@@
    An extensive experimental evaluation of our proposed routing approaches on a real-world testbed, comparing them against an Oracle Router (which has perfect knowledge of the exact number of objects in each image) as well as several baseline techniques, including Random, Round Robin, Lowest Energy Consumption, Lowest Inference Time, and Highest mean Average Precision (mAP). The evaluation leverages multiple realistic datasets, such as the Common Objects in Context (COCO) validation set and a practical surveillance video stream, tested under diverse input scenarios.
@@
The remainder of the paper is organized as follows:Section II discusses the motivation behind this work. Section III presents the system design, problem formulation, and greedy routing approach. Section IV describes the datasets, experimental setup, and results. Section V reviews related literature. Finally, Section VI concludes the paper and outlines directions for future work.
II Motivation
@@
Fig. 2(a) illustrates two contrasting scenes captured by an urban surveillance camera at an intersection. The first image depicts numerous pedestrians crossing the intersection, while the second image shows the same location when it is sparsely populated. In the former high density situation, accurate object detection becomes more critical and at the same time more computationally demanding, especially in safety sensitive applications such as surveillance, traffic monitoring, and crowd management. In contrast, the latter depicts a low-activity scenario in which the computational effort required to achieve reasonable accuracy is reduced, and the risk of critical misdetections is substantially lower. Thus, energy savings can be achieved by using lighter detection models, particularly when edge devices operate under battery or power constraints. This motivates the design of our proposed routing algorithms, which seeks to optimize energy consumption without compromising accuracy where it matters most.
(a) Contrasting pedestrians scenarios: (a) High density scene; (b) Low density scene.
@@
To verify our hypothesis, we conducted a preliminary experiment using the COCO validation dataset, which contains 5,000 images. We categorized the images into two groups based on the number of objects in the image, as determined by the ground truth: one group with a single object per image and another with four or more. We then calculated the mAP and energy consumption per image during inference. As shown in Fig. 2(b)(b), SSD Lite and YOLOv8n exhibit similar accuracy for single-object images, while YOLOv8n significantly outperforms SSD Lite on images with four or more objects, achieving nearly twice better mAP. In terms of energy usage, SSD Lite consumes a consistent amount across both groups, which is approximately 50% lower than YOLOv8n. These results suggest that in less complex scenarios, two different object detection models can achieve similar accuracy, while significant energy savings is possible by using more efficient models such as SSD Lite without compromising accuracy.
(b) Energy Consumption and Accuracy for different group of images across object detection models.
III System Design and Implementation
@@
This section describes the system architecture of ECORE , designed for energy-efficient and performance-aware routing of object detection tasks in an edge environment. Fig. 2 illustrates the overall design. We assume a scenario where a pool of backend edge devices hosts a variety of model–device pairs. Each device runs a dedicated object detection model and exposes an inference API to process images or video frames captured via distributed cameras or embedded sensors.
@@
Cameras or embedded sensors transmit image data or video frames to a central gateway, which performs lightweight preprocessing and analysis. As shown in Fig. 2, the gateway first applies one of the proposed object estimation techniques to infer the number of objects in each incoming image. This count is passed to the routing algorithm (Algorithm 1), which uses it to determine the most appropriate model–device pair.
@@
The routing algorithm begins by identifying the corresponding object count group based on the estimated number of objects, using a predefined set of group rules defined by numeric ranges and labels. It then filters the profiling data to retain only the entries that match the identified group. From this subset, it calculates the maximum achievable mAP and derives a minimum acceptable mAP within a predefined tolerance margin. The data is then filtered again to retain only model–device pairs whose mAP falls within this target range. Among these, the pair with the lowest energy consumption is selected. The gateway forwards the request to the selected endpoint, which processes the image and returns the detected objects and their bounding boxes.
Figure 2: System architecture presentation.
III-A Problem Formulation and Greedy Approach
@@
When running an object detection application in an edge environment, a key challenge lies in routing image requests to the most suitable model-device pair, taking into account trade-offs among accuracy, latency, and energy. Let:
$\mathcal{M}=\{(m_{1},d_{1}),(m_{2},d_{2}),\dots,(m_{n},d_{n})\}$
[…]
@@
$\mathcal{F}=\{i\in\mathcal{M}_{G}\mid\text{mAP}_{i}\geq\text{mAP}_{\text{max}}-\delta_{\text{mAP}}\}$
@@
2:for all $(\texttt{range},\texttt{label})\in\texttt{group\_rules}$ do
3:  if $\texttt{number\_of\_objects}\in\texttt{range}$ then
4:   $\texttt{group}=\texttt{label}$
@@
8:Filter profiling_data to select only rows where $\texttt{object\_count\_group}=\texttt{group}$
@@
11:Compute $\texttt{mAP\_min}=\texttt{max\_mAP}-\texttt{delta\_mAP}$
12:Filter group_data to retain only rows with $\texttt{mAP}\geq\texttt{mAP\_min}$
@@
This greedy algorithm, as detailed in Algorithm 1, provides a principled balance between performance and efficiency, enabling the system to route requests to energy-efficient configurations with only marginal degradation in accuracy. It is particularly suitable for real-time applications deployed in resource constrained environments, where the router must make fast decisions under uncertain and dynamic workloads.
III-B Optimality of the Routing Algorithm
@@
Theorem 1 (Optimality of the Routing Algorithm).
@@
$\mathcal{F}=\left\{i\in\mathcal{M}_{G}\mid\text{mAP}_{i}\geq\text{mAP}_{\max}-\delta_{\text{mAP}}\right\}$
@@
The objective is to minimize energy consumption:
@@
where the feasible set $\mathcal{F}$ includes only model-device pairs in $\mathcal{M}_{G}$ whose accuracy exceeds the threshold $\text{mAP}_{\max}-\delta_{\text{mAP}}$.
@@
As the energy consumption values $e_{i}$ are known from prior uncorrelated profiling, the greedy algorithm that selects
@@
III-C Proposed object count estimation approaches
To avoid negating energy savings, the routing algorithm must remain lightweight, especially its object-counting component. We therefore design and evaluate three approaches for estimating the number of objects in incoming image requests and directing them to the most suitable edge node.”
  1. 1.
    Edge Detection (ED): This approach leverages edge detection which is a fundamental technique in image processing and computer vision used to identify areas of significant intensity or color change, which often correspond to object boundaries. The primary goal is to capture essential structural information. We employ the widely used Canny edge detection algorithm [6], to estimate the number of objects in each image. The resulting object count is then fed into the proposed routing algorithm (Algorithm 1), which selects the optimal model-device pair based on the expected energy–accuracy trade-off governed by $\delta_{\text{mAP}}$. This method is considered lightweight and computationally efficient, though it may yield coarse or approximate estimations.
  2. 2.
@@ III-C Proposed object count estimation approaches
  3. 3.
    Output-Based (OB): This approach leverages temporal continuity in image streams, particularly video, by assuming that consecutive frames contain a similar number of objects. Instead of estimating the object count for every image (frame), the system reuses the count obtained from the output of the previous image (frame), as detected by the last selected model–device pair. The same configuration is retained until a change in object count is observed. The routing algorithm begins with a default estimate (e.g., zero) and forwards the first image to an edge node. Once the detection results are returned, the router extracts the actual number of objects and uses it for subsequent routing decisions. If the object count remains stable, the system continues using the same model–device pair; if a change is detected, the routing decision is updated accordingly. This adaptive strategy avoids per-image (frame) estimation, reducing computational overhead and improving energy efficiency.
IV Performance Evaluation
IV-A Experimental Setup
IV-A1 Dataset
In this study, we utilize three distinct datasets, two of which are derived from the COCO validation set [7] and video stream frames, to evaluate the routing algorithm. The first is the Original COCO Validation Dataset, consisting of 5,000 images annotated with ground truth labels, including object classes and bounding boxes spanning 80 object categories. As this dataset reflects natural scene variability, the number of objects per image varies significantly. Fig. 3 illustrates the distribution of object counts, highlighting this diversity.